# On Convolutions, Intrinsic Dimension, and Diffusion Models

### v1

@@ Abstract
Generative models aim to learn the distribution $p$ that gave rise to observed data, and when the manifold hypothesis holds, $p$ has low-dimensional support. Research studying generative models through the lens of the manifold hypothesis has recently proliferated (Dai & Wipf, 2019; Zhang et al., 2020; Brehmer & Cranmer, 2020; Kim et al., 2020; Arbel et al., 2021; Caterini et al., 2021; Horvat & Pfister, 2021; Ross & Cresswell, 2021; Cunningham et al., 2022; Loaiza-Ganem et al., 2022a; b; Ross et al., 2023; Sorrenson et al., 2024; Loaiza-Ganem et al., 2024), and an important finding from this line of research is that diffusion models (DMs; Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) – which are state-of-the-art generative models – can successfully learn distributions with low-dimensional support (Pidstrigach, 2022; De Bortoli, 2022). When a generative model, such as a DM, correctly learns $p$ it must therefore learn the corresponding intrinsic dimensions of its support as well; indeed, various works have shown that LID estimators can be extracted from generative models and that these new estimators avoid the aforementioned pitfalls of traditional ones (Tempczyk et al., 2022; Zheng et al., 2022; Horvat & Pfister, 2024; Stanczuk et al., 2024; Kamkari et al., 2024b).
@@
Throughout our work, we will follow the definition of a $d$-dimensional manifold as being locally homeomorphic to $\mathbb{R}^{d}$ (Lee, 2012), meaning that the dimension of a manifold does not vary within the manifold. $\mathcal{M}$ will denote, depending on context, either an embedded $d$-dimensional submanifold of $\mathbb{R}^{D}$, or a countable disjoint union of embedded submanifolds of potentially varying dimensions. We will assume that data of interest is supported on $\mathcal{M}$ and generated according to a probability distribution admitting a density $p:\mathcal{M}\rightarrow\mathbb{R}$ [^1] Note that $p$ is not a density with respect to the Lebesgue measure on $\mathbb{R}^{D}$, but rather with respect to the Riemannian measure (or volume form if it exists) on $\mathcal{M}$; see the work of Loaiza-Ganem et al. (2024) for a discussion about this setup. Writing $\mathcal{M}$ as $\mathcal{M}=\cup_{j}\mathcal{M}_{j}$ where $\mathcal{M}_{j}$ is $d_{j}$-dimensional and $\mathcal{M}_{i}\cap\mathcal{M}_{j}=\emptyset$ for $i\neq j$, the local intrinsic dimension of $x\in\mathcal{M}$ is defined as the dimension of the manifold that $x$ belongs to, i.e. $\texttt{LID}(x)=d_{i}$ if $x\in\mathcal{M}_{i}$.
We will make heavy use of isotropic Gaussian densities with log standard deviation $\delta\in\mathbb{R}$, and for a mean $\mu\in\mathbb{R}^{D}$, we will denote the corresponding density evaluated at $x\in\mathbb{R}^{D}$ as
@@
[…]
@@
where $\alpha:\mathbb{R}^{D}\times[0,1]\rightarrow\mathbb{R}^{D}$ and $\beta:[0,1]\rightarrow\mathbb{R}$ are fixed functions which are specified as hyperparameters, and $W_{t}$ is a $D$-dimensional Brownian motion. We will denote the density of $X_{t}$ implied by Equation 10 as $p(\,\cdot\,,t)$, and note that $p(\,\cdot\,,0)=p$ [^3] Note that $p(\,\cdot\,,0)$ and $p(\,\cdot\,,t)$ for $t>0$ are not densities in the same sense; only the latter are Lebesgue densities. Here $t\in[0,1]$ corresponds to an artificial time variable specifying how much noise has been added to the data through the SDE, with $t=0$ corresponding to no noise, and with $t=1$ corresponding to $p(\,\cdot\,,1)$ being “almost pure noise”. A surprising result by Anderson (1982) and Haussmann & Pardoux (1986) shows that the reverse process, $Y_{t}\coloneqq X_{1-t}$, also obeys a (backward) SDE,
@@
The function $\alpha$ is often chosen as $\alpha(x,t)=\gamma(t)x$ for some function $\gamma:[0,1]\rightarrow\mathbb{R}$. Under this choice, the transition kernel corresponding to Equation 10 is given by (Särkkä & Solin, 2019)
@@
The correctness of FLIPD as an estimator of LID therefore hinges on Equation 1 holding in general, and as previously mentioned, Kamkari et al. (2024b) only proved Equation 1 in the case where $\mathcal{M}$ is an affine submanifold of $\mathbb{R}^{D}$.
@@
By adding different levels of noise $\delta$ to data and training a normalizing flow for each noise level, Tempczyk et al. (2022) estimate $\log\varrho_{\mathcal{N}}(x,\delta)$ for various values of $\delta$; they then fit a linear regression predicting these estimates from $\delta$ – the resulting slope is an estimate $\texttt{LID}(x)-D$ thanks to Equation 17. When proposing FLIPD, Kamkari et al. (2024b) noted that Equation 17 provides informal intuition for why Equation 1 holds: if the $\mathcal{O}(1)$ term was constant with respect to $\delta$, then Equation 17 would imply Equation 1. In this sense, proving Equation 1 boils down to showing that the $\mathcal{O}(1)$ term indeed behaves like a constant as $\delta\rightarrow-\infty$, or more formally, that the limit of its derivative with respect to $\delta$ converges to 00.
@@
[…]
@@
As previously discussed, Corollary 2 immediately links $\texttt{LID}(x)$ and $\mathbb{P}_{X\sim p}(X\in B_{D}(x,e^{\delta}))$ through Equation 9. Albeit Corollary 1 is more relevant due to its connection to DMs and FLIPD, Corollary 2 still succeeds at connecting two quantities of interest in machine learning. Some older estimators of LID use a given a dataset sampled from $p$ by regressing the log proportion of datapoints in $B_{D}(x,e^{\delta})$ against $\delta$, and then estimating $\texttt{LID}(x)$ as the corresponding slope (Pettis et al., 1979; Grassberger & Procaccia, 1983). These works provided an intuitively correct but not rigorous mathematical justification behind this estimator, which Equation 9 rigorously justifies.
@@
The bound in Equation 38 decreases monotonically to 00 as $\delta\rightarrow-\infty$, and the bound in Equation 39 does so as well provided that $\delta<\log\xi-\tfrac{1}{2}\log(D+2)$ (which makes the derivative of the log of the bound with respect to $\delta$ positive). It follows that $\mathcal{N}_{D}(x-x^{\prime};0,\delta)$ and $\|x-x^{\prime}\|^{2}e^{-2\delta}\mathcal{N}_{D}(x-x^{\prime};0,\delta)$ are both uniformly bounded over $x^{\prime}$ and $\delta$ when $x^{\prime}\in\mathcal{M}_{j}$ for $j\neq i$ and $\delta<\min(\tfrac{1}{2}\log\tfrac{\xi}{2},\log\xi-\tfrac{1}{2}\log(D+2))$. The dominated convergence theorem then guarantees that, for $j\neq i$:
@@
being violated. Indeed, Kamkari et al. (2024b) reported that in some instances, FLIPD can produce negative estimates of LID. Since Kamkari et al. (2024b) only proved Equation 1 when $\mathcal{M}$ is an affine submanifold of $\mathbb{R}^{D}$, the possibility remained that this negativity was due to the result not holding more generally; a consequence of Corollary 1 is that negative estimates can only be caused by $\hat{s}$ having imperfectly learned $s$. We believe that quantifying how much $\texttt{FLIPD}\left(x;\delta_{0}\right)$ differs from $\texttt{LID}(x)$ as a function of the error incurred by $\hat{s}$ to approximate $s$ is another interesting direction for future research.
@@ Appendix A Proofs for the Gaussian Case
[…]
@@
Without loss of generality, assume $\delta<\log{r}-\tfrac{1}{2}\log{(d+2)}$, then the integrand will decrease as $\delta$ decreases (the derivative of the $\log$ of the integrand with respect to $\delta$ is positive) and it converges pointwise to 00. It follows that the integrand is bounded for negative enough $\delta$, and the result follows from the dominated convergence theorem. ∎
@@
Let $W$ be an open and bounded subset of $\mathbb{R}^{d}$ with $0\in W$. Let $f:W\rightarrow\mathbb{R}$ continuous at 00 and bounded on $W$. Then we have
@@ Appendix B Proofs for the Uniform Case
There exists an open set $W$ containing 00 such that $\nabla G$ is well-defined and non-vanishing on $W\setminus\{0\}$.
@@
[…]
@@
Since $u(0)=0$, we have $u_{k}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $u_{k}(0)=0$. Thus, by Claim 4, $u_{k}(y)=\sum_{j}g_{kj}(y)y_{j}$ for some smooth $g_{kj}$. Since all the derivatives of $u$ vanish at 00, for any $k,l$, we have $0=u_{k,l}(0)=g_{kl}(0)+\sum_{j}g_{kj,l}(0)\cdot 0=g_{kl}(0)$. Thus, once again by Claim 4, $g_{kl}(y)=\sum_{j}h_{klj}(y)y_{j}$ for smooth $h_{klj}$. Thus we have $u_{k}(y)=\sum_{i}g_{ki}(y)y_{i}=\sum_{i,j}h_{kij}(y)y_{i}y_{j}$. Thus $(A_{ij})_{k}=h_{kij}$, which completes the proof. ∎
@@
Since, by Claim 5, both $\frac{2y^{\top}(Du)u}{y^{\top}y}$ and $\frac{u^{\top}(Du)^{\top}(Du)u}{y^{\top}y}$ approach 00 as $y\rightarrow 0$, there is an open set around $y=0$ such that $\|y+(Du)u\|^{2}\neq 0$ in this open set when $y\neq 0$. This means that there exists an open set $W$ around $y=0$ such that $\nabla G$ is well-defined and non-vanishing on $W\setminus\{0\}$. ∎
@@
$\psi(y)=\left\{\begin{array}[]{rcl}1&\mbox{ if }&\delta_{a}\leq G(y)\leq\delta_{b}\\ 0&\mbox{ if }&G(y)<\delta_{a^{\prime}}\mbox{ or }G(y)>\delta_{b^{\prime}}\\ \end{array}\right.$ (133)
@@
It suffices to calculate div $\mbox{div }\frac{\nabla G}{\|\nabla G\|^{2}}=\mbox{div }H(y)(y+(Du)u)$. Using Einstein summation notation, we have
@@
(div $\left(\mbox{div }\frac{\nabla G}{\|\nabla G\|^{2}}\right)(0)=H(0)(d+0+0)+H_{i}(0)(0+0)=d.$ (143)
Thus for $\epsilon>0$, there exists $\kappa^{\prime}>0$ such that if $\|y\|<\kappa^{\prime}$, then |(div $\left|(\mbox{div }\frac{\nabla G}{\|\nabla G\|^{2}})(y)-d\right|<\epsilon$. Choose $K^{\prime\prime}$ such that $V_{\delta}\subset B_{d}(0,\kappa^{\prime})$ for $\delta<K^{\prime\prime}$. Then

### v2

@@ Abstract
Generative models aim to learn the distribution $p$ that gave rise to observed data, and when the manifold hypothesis holds, $p$ has low-dimensional support. Research studying generative models through the lens of the manifold hypothesis has recently proliferated (Dai & Wipf, 2019; Zhang et al., 2020; Brehmer & Cranmer, 2020; Kim et al., 2020; Arbel et al., 2021; Caterini et al., 2021; Horvat & Pfister, 2021; Ross & Cresswell, 2021; Cunningham et al., 2022; Loaiza-Ganem et al., 2022a; b; Ross et al., 2023; Sorrenson et al., 2024; Loaiza-Ganem et al., 2024; Ventura et al., 2025), and an important finding from this line of research is that diffusion models (DMs; Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) – which are state-of-the-art generative models – can successfully learn distributions with low-dimensional support (Pidstrigach, 2022; De Bortoli, 2022). When a generative model, such as a DM, correctly learns $p$ it must therefore learn the corresponding intrinsic dimensions of its support as well; indeed, various works have shown that LID estimators can be extracted from generative models and that these new estimators avoid the aforementioned pitfalls of traditional ones (Tempczyk et al., 2022; Zheng et al., 2022; Horvat & Pfister, 2024; Stanczuk et al., 2024; Kamkari et al., 2024b).
@@
Throughout our work, we will follow the definition of a $d$-dimensional manifold as being locally homeomorphic to $\mathbb{R}^{d}$ (Lee, 2012), meaning that the dimension of a manifold does not vary within the manifold. $\mathcal{M}$ will denote, depending on context, either an embedded $d$-dimensional submanifold of $\mathbb{R}^{D}$, or a countable disjoint union of embedded submanifolds of potentially varying dimensions. We will assume that data of interest is supported on $\mathcal{M}$ and generated according to a probability distribution admitting a density $p:\mathcal{M}\rightarrow\mathbb{R}$. [^1] Note that $p$ is not a density with respect to the Lebesgue measure on $\mathbb{R}^{D}$, but rather with respect to the Riemannian measure (or volume form if it exists) on $\mathcal{M}$; see the work of Loaiza-Ganem et al. (2024) for a discussion about this setup. Writing $\mathcal{M}$ as $\mathcal{M}=\cup_{j}\mathcal{M}_{j}$ where $\mathcal{M}_{j}$ is $d_{j}$-dimensional and $\mathcal{M}_{i}\cap\mathcal{M}_{j}=\emptyset$ for $i\neq j$, the local intrinsic dimension of $x\in\mathcal{M}$ is defined as the dimension of the manifold that $x$ belongs to, i.e. $\texttt{LID}(x)=d_{i}$ if $x\in\mathcal{M}_{i}$.
We will make heavy use of isotropic Gaussian densities with log standard deviation $\delta\in\mathbb{R}$, i.e. with covariance matrix $e^{2\delta}I_{D}$; for a mean $\mu\in\mathbb{R}^{D}$, we will denote the corresponding density evaluated at $x\in\mathbb{R}^{D}$ as
@@
and where $\|\cdot\|$ will always refer to the Euclidean norm. We will sometimes find it convenient to write $\mathcal{N}_{d}(x;\mu,\delta)$ where $d<D$ and $x,\mu\in\mathbb{R}^{D}$, with the understanding that
@@
where $\alpha:\mathbb{R}^{D}\times[0,1]\rightarrow\mathbb{R}^{D}$ and $\beta:[0,1]\rightarrow\mathbb{R}$ are fixed functions which are specified as hyperparameters, and $W_{t}$ is a $D$-dimensional Brownian motion. We will denote the density of $X_{t}$ implied by Equation 10 as $p(\,\cdot\,,t)$, and note that $p(\,\cdot\,,0)=p$. [^3] Note that $p(\,\cdot\,,0)$ and $p(\,\cdot\,,t)$ for $t>0$ are not densities in the same sense; only the latter are Lebesgue densities. Here $t\in[0,1]$ corresponds to an artificial time variable specifying how much noise has been added to the data through the SDE, with $t=0$ corresponding to no noise, and with $t=1$ corresponding to $p(\,\cdot\,,1)$ being “almost pure noise”. A surprising result by Anderson (1982) and Haussmann & Pardoux (1986) shows that the reverse process, $Y_{t}\coloneqq X_{1-t}$, also obeys a (backward) SDE,
@@
The function $\alpha$ is often chosen as $\alpha(x,t)=\gamma(t)x$ for some function $\gamma:[0,1]\rightarrow\mathbb{R}$. Under this choice, the transition kernel corresponding to Equation 10 is known (Särkkä & Solin, 2019), and it is given by
@@
The correctness of FLIPD as an estimator of LID therefore hinges on Equation 1 holding in general, and as previously mentioned, Kamkari et al. (2024b) only proved Equation 1 in the case where $\mathcal{M}$ is an affine submanifold of $\mathbb{R}^{D}$. We refer the reader to Kamkari et al. (2024b) for a more thorough derivation of FLIPD, along with illustrative examples, and empirical results.
@@
By adding different levels of noise $\delta$ to data and training a normalizing flow for each noise level, Tempczyk et al. (2022) estimate $\log\varrho_{\mathcal{N}}(x,\delta)$ for various values of $\delta$; they then fit a linear regression predicting these estimates from $\delta$ – the resulting slope is an estimate $\texttt{LID}(x)-D$ thanks to Equation 17. When proposing FLIPD, Kamkari et al. (2024b) noted that Equation 17 provides informal intuition for why Equation 1 holds: if the $\mathcal{O}(1)$ term was constant with respect to $\delta$, then Equation 17 would imply Equation 1. In this sense, proving Equation 1 boils down to showing that the $\mathcal{O}(1)$ term indeed behaves like a constant as $\delta\rightarrow$-$\infty$, or more formally, that the limit of its derivative with respect to $\delta$ converges to 0.
Several works have studied the interplay between DMs and the manifold hypothesis in contexts beyond LID estimation: Pidstrigach (2022) first showed that if the error between $\hat{s}$ and $s$ is bounded, then DMs recover the correct support, thus showing that DMs can learn manifolds. De Bortoli (2022) refined this result with error bounds in Wasserstein distance, and various follow-up works have provided finite-sample estimation rates (Chen et al., 2023; Oko et al., 2023; Tang & Yang, 2024; Wang et al., 2024). Lastly, Chen et al. (2024) observed that, under the manifold hypothesis, the Jacobian of the score function has low rank, and they further leveraged this observation for image editing.
@@
Let $\mathcal{M}=\cup_{j}\mathcal{M}_{j}$, where $\mathcal{M}_{j}$ is a smooth $d_{j}$-dimensional embedded submanifold of $\mathbb{R}^{D}$ with $\xi\coloneqq\min_{i\neq j}\inf_{x_{i}\in\mathcal{M}_{i},x_{j}\in\mathcal{M}_{j}}\|x_{i}-x_{j}\|>0$. Let $p$ be a probability density function on $\mathcal{M}$, which we write as $p(x)=\pi_{j}p_{j}(x)$ for $x\in\mathcal{M}_{j}$, where for every $j$, $p_{j}$ is a probability density on $\mathcal{M}_{j}$ and $\pi_{j}>0$, and $\sum_{j}\pi_{j}=1$. [^4] Note that, because $\xi>0$, without loss of generality any density $p$ on $\mathcal{M}$ can be written as $p(x)=\pi_{j}p_{j}(x)$ when $x\in\mathcal{M}_{j}$. Let $x\in\mathcal{M}_{i}$ for some $i$ be such that $p_{i}$ is continuous at $x$, $p_{i}(x)>0$, and $\int_{\mathcal{M}_{i}}p_{i}(x^{\prime})\texttt{dist}_{\mathcal{M}_{i}}^{2}(x,x^{\prime}){\textnormal{d}}x^{\prime}<\infty$, where $\texttt{dist}_{\mathcal{M}_{i}}$ denotes the geodesic distance on $\mathcal{M}_{i}$ obtained from the induced Riemannian metric on $\mathcal{M}_{i}$. Then,
@@
As previously discussed, Corollary 2 immediately links $\texttt{LID}(x)$ and $\mathbb{P}_{X\sim p}(X\in B_{D}(x,e^{\delta}))$ through Equation 9. Albeit Corollary 1 is more relevant due to its connection to DMs and FLIPD, Corollary 2 still succeeds at connecting two quantities of interest in machine learning despite lacking these connections. Some older estimators of LID use a given a dataset sampled from $p$ by regressing the log proportion of datapoints in $B_{D}(x,e^{\delta})$ against $\delta$, and then estimating $\texttt{LID}(x)$ as the corresponding slope (Pettis et al., 1979; Grassberger & Procaccia, 1983). These works provided an intuitively correct but not rigorous mathematical justification behind this estimator, which Equation 9 rigorously justifies.
@@
The bound in Equation 38 decreases monotonically to 0 as $\delta\rightarrow$-$\infty$, and the bound in Equation 39 does so as well provided that $\delta<\log\xi-\tfrac{1}{2}\log(D+2)$ (which makes the derivative of the log of the bound with respect to $\delta$ positive). It follows that $\mathcal{N}_{D}(x-x^{\prime};0,\delta)$ and $\|x-x^{\prime}\|^{2}e^{-2\delta}\mathcal{N}_{D}(x-x^{\prime};0,\delta)$ are both uniformly bounded over $x^{\prime}$ and $\delta$ when $x^{\prime}\in\mathcal{M}_{j}$ for $j\neq i$ and $\delta<\min(\tfrac{1}{2}\log\tfrac{\xi}{2},\log\xi-\tfrac{1}{2}\log(D+2))$. The dominated convergence theorem then guarantees that, for $j\neq i$:
@@
being violated. Indeed, Kamkari et al. (2024b) reported that in some instances, FLIPD can produce negative estimates of LID. Since Kamkari et al. (2024b) only proved Equation 1 when $\mathcal{M}$ is an affine submanifold of $\mathbb{R}^{D}$, the possibility remained that this negativity was due to the result not holding more generally; a consequence of Corollary 1 is that negative estimates can only be caused by $\hat{s}$ having imperfectly learned $s$. We believe that quantifying how much $\texttt{FLIPD}\left(x;\delta_{0}\right)$ differs from $\texttt{LID}(x)$ as a function of the error incurred by $\hat{s}$ to approximate $s$ is another relevant problem. In particular, note that if we had bounds on both the error between $s$ and $\hat{s}$, and on the error between $\nabla s$ and $\nabla\hat{s}$, as $\delta\rightarrow$-$\infty$, we could trivially obtain a bound on $|\nu(s,x,t(\delta))-\nu(\hat{s},x,t(\delta))|$ (see Equation 15); in turn, this would provide a bound on the LID estimation error. While we are aware of learning theory work bounding the error between $s$ and $\hat{s}$ (Chen et al., 2023; Oko et al., 2023; Tang & Yang, 2024; Wang et al., 2024), we are unaware of any existing bound on the error between $\nabla s$ and $\nabla\hat{s}$; finding such a bound is another interesting direction for future work.
@@ Appendix A Proofs for the Gaussian Case
To compute this limit, without loss of generality, we perform a translation so that $x=0$. Note that this translation preserves distance. Then perform an orthogonal transformation on $\mathbb{R}^{D}$ such that the tangent plane of $\mathcal{M}$ at $x=0$ is the $d$-subspace spanned by the first $d$ coordinate vectors. Denote these coordinates by $(y_{1},\ldots,y_{d})$. As $\mathcal{M}$ is smooth, $\mathcal{M}$ can be parametrized by $\{y,u(y)\}$ near $x=0$, where $y=(y_{1},\ldots,y_{d})$ and $u$ is a function from some open $U\subset\mathbb{R}^{d}$ to $\mathbb{R}^{D-d}$, such that $u(0)=0$ and (Du)(0)=0(Du)(0)=0.555To avoid using non-standard notation for derivatives, we overload notation and use $D$ to refer to both the derivative operator and the ambient dimension. The meaning of $D$ will always be clear from context. In other words, $(y_{1},\ldots,y_{d})$ is a local coordinate system of $\mathcal{M}$ on $U$, with the induced Riemannian metric $g$. Note that $g=I_{d}$ at $x=0$ as $(Du)(0)=0$. This means that $U$ is diffeomorphic to an open set in $\mathcal{M}$ via $\phi(x)=(x,u(x))$. Shrink $U$ if necessary such that $U$ is bounded, i.e. $R\coloneqq\sup_{x^{\prime}\in\phi(U)}\|x^{\prime}\|<\infty$. Further shrink $U$ if necessary such that $B_{D}(0,R)\cap\mathcal{M}=\phi(U)$. In other words, $x^{\prime}\in\mathcal{M},\|x^{\prime}\|<R\Longleftrightarrow x^{\prime}\in\phi(U)$.
@@
Without loss of generality, assume $\delta<\log{r}-\tfrac{1}{2}\log{(d+2)}$, then the integrand will decrease as $\delta$ decreases (the derivative of the $\log$ of the integrand with respect to $\delta$ is positive) and it converges pointwise to 0. It follows that the integrand is bounded for negative enough $\delta$, and the result follows from the dominated convergence theorem. ∎
@@
Let $W$ be an open and bounded subset of $\mathbb{R}^{d}$ with $0\in W$. Let $f:W\rightarrow\mathbb{R}$ continuous at 0 and bounded on $W$. Then we have
@@ Appendix B Proofs for the Uniform Case
There exists an open set $W$ containing 0 such that $\nabla G$ is well-defined and non-vanishing on $W\setminus\{0\}$.
@@
$f(y)=f(y)-f(0)=\int_{0}^{1}\frac{\partial}{\partial t}f(ty)dt=\int_{0}^{1}\sum_{i}y_{i}\frac{\partial f}{\partial y_{i}}f(ty)dt.$ (127)
Thus $g_{i}=\int_{0}^{1}\frac{\partial f}{\partial y_{i}}f(ty)dt$ and $g_{i}$ is smooth. ∎
@@
Since $u(0)=0$, we have $u_{k}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $u_{k}(0)=0$. Thus, by Claim 4, $u_{k}(y)=\sum_{j}g_{kj}(y)y_{j}$ for some smooth $g_{kj}$. Since all the derivatives of $u$ vanish at 0, for any $k,l$, we have $0=u_{k,l}(0)=g_{kl}(0)+\sum_{j}g_{kj,l}(0)\cdot 0=g_{kl}(0)$. Thus, once again by Claim 4, $g_{kl}(y)=\sum_{j}h_{klj}(y)y_{j}$ for smooth $h_{klj}$. Thus we have $u_{k}(y)=\sum_{i}g_{ki}(y)y_{i}=\sum_{i,j}h_{kij}(y)y_{i}y_{j}$. Thus $(A_{ij})_{k}=h_{kij}$, which completes the proof. ∎
@@
Since, by Claim 5, both $\frac{2y^{\top}(Du)u}{y^{\top}y}$ and $\frac{u^{\top}(Du)^{\top}(Du)u}{y^{\top}y}$ approach 0 as $y\rightarrow 0$, there is an open set around $y=0$ such that $\|y+(Du)u\|^{2}\neq 0$ in this open set when $y\neq 0$. This means that there exists an open set $W$ around $y=0$ such that $\nabla G$ is well-defined and non-vanishing on $W\setminus\{0\}$. ∎
@@
[…]
@@
It suffices to calculate $\mbox{div }\frac{\nabla G}{\|\nabla G\|^{2}}=\mbox{div }H(y)(y+(Du)u)$. Using Einstein summation notation, we have
@@
$\left(\mbox{div }\frac{\nabla G}{\|\nabla G\|^{2}}\right)(0)=H(0)(d+0+0)+H_{i}(0)(0+0)=d.$ (143)
Thus for $\epsilon>0$, there exists $\kappa^{\prime}>0$ such that if $\|y\|<\kappa^{\prime}$, then $\left|(\mbox{div }\frac{\nabla G}{\|\nabla G\|^{2}})(y)-d\right|<\epsilon$. Choose $K^{\prime\prime}$ such that $V_{\delta}\subset B_{d}(0,\kappa^{\prime})$ for $\delta<K^{\prime\prime}$. Then